# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v2

@@ 4. Checking the existence of PR causes and the PR conditions
We now turn to algorithms for checking whether a given set $\mathsf{Cause}$ is an SPR or GPR cause for $\mathsf{Eff}$ in $\mathcal{M}$. Since the minimality condition (for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$) of PR causes is verifiable by standard model checking techniques in polynomial time, we concentrate on checking the probability-raising conditions (S) and (G). For Markov chains, both (SPR) and (GPR) can be checked in polynomial time by computing the corresponding probabilities. So, the interesting case is checking the PR conditions in MDPs. In case of SPR causality this is closely related to the existence of PR causes and solvable in polynomial time (Section 4.1), while checking the GPR condition is more complex and polynomially reducible to (the non-solvability of) a quadratic constraint system (Section 4.2).
We start by stating that for (S) and (G), it suffices to consider kind of worst-case scheduler, which are minimizing the probability to reach an effect state from every cause state. For this we transform the MDP in question according to the cause candidate in question.
@@
{exa}
@@
Using
$\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff})$ $=$ $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\psi_{c}\wedge\Diamond\mathsf{Eff})\ +\ \mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}((\neg\psi_{c})\wedge\Diamond\mathsf{Eff}),$
we can equivalently convert condition (3) for $c$ and $\mathfrak{S}$ to
@@
Suppose $\mathsf{Cause}$ is an GPR cause for $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$. Then, $\mathsf{Cause}$ is an GPR cause for $\mathsf{Eff}$ in $\mathcal{M}$.
@@
From the assumption that $\mathsf{Cause}$ is an GPR cause for $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$, we can conclude that the GPR condition (GPR) holds for all schedulers $\mathfrak{S}$ that satisfy
@@
Pick an MD-scheduler $\mathfrak{T}$ that minimizes the probability to reach $\mathsf{Eff}$ from every state. In particular, $w_{c}=w_{\pi}^{\mathfrak{T}}\leqslant w_{\pi}^{\mathfrak{S}}$ for every state $c\in\mathsf{Cause}$ and every path $\pi\in\Pi_{c}$ (recall that $w_{c}=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$). Moreover, the scheduler $\mathfrak{S}$ can be transformed into a scheduler $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0,\min}$ that is “equivalent” to $\mathfrak{S}$ with respect to the global probability-raising condition. More concretely, let $\mathfrak{S}_{\mathfrak{T}}$ denote the scheduler that behaves as $\mathfrak{S}$ as long as $\mathfrak{S}$ has not yet visited a state in $\mathsf{Cause}$ and behaves as $\mathfrak{T}$ as soon as a state in $\mathsf{Cause}$ has been reached. Thus, $p_{\pi}^{\mathfrak{S}}=p_{\pi}^{\mathfrak{S}_{\mathfrak{T}}}$ and $\mathit{res}(\mathfrak{S}_{\mathfrak{T}},\pi)=\mathfrak{T}$ for each $\pi\in\Pi_{c}$. This yields that the probability to reach $c\in\mathsf{Cause}$ from $\mathsf{init}$ is the same under $\mathfrak{S}$ and $\mathfrak{S}_{\mathfrak{T}}$, i.e., $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond c)=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\Diamond c)$. Therefore $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Cause})=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\Diamond\mathsf{Cause})$. The latter implies that $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0}$, and hence $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0,\min}$. Moreover, $\mathfrak{S}$ and $\mathfrak{S}_{\mathfrak{T}}$ reach $\mathsf{Eff}$ without visiting $\mathsf{Cause}$ with the same probability, i.e., $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,\mathsf{Eff})=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,\mathsf{Eff})$.
But this yields: if (5) holds for $\mathfrak{S}_{\mathfrak{T}}$ then (5) holds for $\mathfrak{S}$. As (5) holds for $\mathfrak{S}_{\mathfrak{T}}$ by assumption, this completes the proof.
@@ 4.1. Checking the strict probability-raising condition and the existence of causes
First, we show the soundness of Algorithm 4.1. By the virtue of Lemma 10 stating the soundness of the tranformation $\mathcal{M}$ to $\mathcal{M}_{[c]}$ it suffices to show that Algorithm 4.1 returns the correct answers “yes” or “no” when the task is to check whether the singleton $\mathsf{Cause}=\{c\}$ is an SPR cause in $\mathcal{N}=\mathcal{M}_{[c]}$. Recall the notation $q_{s}=\mathrm{Pr}^{\max}_{\mathcal{M}_{[c]},s}(\Diamond\mathsf{Eff})$. We abbreviate $q=q_{\mathsf{init}}$. Note that $(\neg\mathsf{Cause})\,\mathrm{U}\,c$ is equivalent to $\Diamond c$ as $c\in\mathsf{Cause}$.
@@
    of Algorithm 4.1 (i.e., if $q<w_{c}$), the answer “yes” is sound as then $\mathrm{Pr}^{\max}_{\mathcal{N}}(\Diamond\mathsf{Eff})=q<w_{c}$.
@@
    Pick an MD-scheduler $\mathfrak{S}$ of $\mathcal{M}^{\max}_{[c]}$ such that $c$ is reachable from $\mathsf{init}$ via $\mathfrak{S}$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N},s}(\Diamond\mathsf{Eff})=q_{s}$ for all states $s$. Hence, (SPR) does not hold for $c$ and the scheduler $\mathfrak{S}$.
@@
    We have the property that $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond c)=0$ for all schedulers $\mathfrak{S}$ for $\mathcal{N}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})=q=w_{c}$. But then $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond c)>0$ implies $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})<w_{c}$ as required in (SPR).
@@
    there is a state $c_{0}\in S\setminus\mathsf{Eff}$ such that the singleton $\{c_{0}\}$ is an SPR cause (and therefore a GRP cause) for $\mathsf{Eff}$ in $\mathcal{M}$.
@@ 4.2. Checking the global probability-raising condition
The terminal states $\mathsf{eff}_{\mathsf{unc}}$, $\mathsf{eff}_{\mathsf{cov}}$, $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{noeff}_{\mathsf{tn}}$ are pairwise distinct. $\mathcal{M}$ can have further terminal states representing true negatives. These can be identified with $\mathsf{noeff}_{\mathsf{tn}}$.
@@
We justify the assumptions as we can transform $\mathcal{M}$ into a new MDP of the same asymptotic size satisfying the above assumptions. Thanks to Lemma 10, we may suppose that $\mathcal{M}=\mathcal{M}_{[\mathsf{Cause}]}$ without changing the satisfaction of (G). Thus, from cause states $c\in\mathsf{Cause}$ there are only two outgoing transitions, either to a terminal effect state $\mathsf{eff}$ with probability $\mathrm{Pr}^{\min}_{c}(\lozenge\mathsf{Eff})$ or to a terminal non-effect state $\mathsf{noeff}$ with the remaining probability (see Notation 4). We then may rename the effect state $\mathsf{eff}$ and the non-effect state $\mathsf{noeff}$ reachable from $\mathsf{Cause}$ into $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$, respectively. Furthermore, we collapse all other effect states into a single state $\mathsf{eff}_{\mathsf{unc}}$ and all true negative states into $\mathsf{noeff}_{\mathsf{tn}}$. Similarly, by renaming and possibly duplicating terminal states we also suppose that $\mathsf{noeff}_{\mathsf{fp}}$ has no other incoming transitions than the $\gamma$-transitions from the states in $\mathsf{Cause}$. This ensures (A1) and (A2). For (A3) consider the set $T$ of terminal states in the MDP obtained so far. We remove all non-trivial end components by switching to the MEC-quotient [dA97], i.e., we collapse all states that belong to the same MEC $\mathcal{E}$ into a single state $s_{\mathcal{E}}$ while ignoring the actions inside $\mathcal{E}$. Additionally, we add a fresh $\tau$-transition from the states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ (i.e., $P(s_{\mathcal{E}},\tau,\mathsf{noeff}_{\mathsf{tn}})=1$). The $\tau$-transitions from states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ mimic cases where the scheduler of the original MDP enters the end component $\mathcal{E}$ and stays there forever.
@@
{exa}
For a demonstration of the above described transformations consider the abstract MDP $\mathcal{M}$ from Figure 12, where the dotted circles correspond to sets of states in the MDP. The MDP already has the form $\mathcal{M}=\mathcal{M}_{[c]}$. Now rename the effect state $\mathsf{eff}$ reachable from $\mathsf{Cause}$ to $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}$ to $\mathsf{noeff}_{\mathsf{fp}}$. All effect states which are not reachable from $\mathsf{Cause}$ are collapsed into the state $\mathsf{eff}_{\mathsf{unc}}$. In this example there are no terminal non-effect states which are not reachable from $c$. These would be collapsed to $\mathsf{noeff}_{\mathsf{tn}}$. Taking the MEC quotient collapses the MECs to states $s_{\mathcal{E}_{i}}$ while keeping outgoing transitions. Additionally, there is a fresh outgoing action $\tau$ from the states $s_{\mathcal{E}_{i}}$ to the fresh state $\mathsf{noeff}_{\mathsf{tn}}$. Thus we get $\mathcal{N}$ from Figure 12.
@@
Figure 11. Example MDP $\mathcal{M}$ not satisfying
@@
The violation of (GPR-1) in Lemma 21 can be reformulated in terms of the frequency-variables as follows where $x_{\mathsf{Cause}}$ is an abbreviation for $\sum_{c\in\mathsf{Cause}}x_{c}$:
@@
Finally, we can reformulate the condition $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ by frequency variables:
@@
We can now prove that deciding the GPR condition can be done in $\mathrm{coNP}$, Theorem 17.
@@
There are solutions to (S1)-(S3) and (S5) that violate (S4). The space of feasible points for conditions (S1)-(S3) and (S5) is connected. Furthermore, the right hand side of (S4)
@@
where the variables $y$ are as the original variables $x$ after the variables in $X^{\prime}$ have been removed. Now, we apply the implicit function theorem: Observe that
@@
The model transformation we used for assumptions (A1)-(A3), however, does affect the memory requirements of scheduler. We may further restrict the MR-schedulers necessary to witness non-causality under assumptions (A1)-(A3). For the following lemma, recall that $\tau$ is the action of the MEC quotient used for the extra transition from states representing MECs to a new trap state (see also assumption (A3)).
@@
To explain how to derive $\mathfrak{T}_{i}$ from $\mathfrak{T}_{i-1}$, let $i\in\{1,\ldots,k\}$, $\mathfrak{V}=\mathfrak{T}_{i-1}$, $u=u_{i}$ and $y=1{-}\mathfrak{V}(u)(\tau)$. Then, $0<y<1$ (as $u\in U$ and by definition of $U$) and $y=\sum_{\alpha\in\mathit{Act}(u)\setminus\{\tau\}}\mathfrak{V}(u)(\alpha)$. For $x\in[0,1]$, let $\mathfrak{V}_{x}$ denote the MR-scheduler that agrees with $\mathfrak{V}$ for all states but $u$, for which $\mathfrak{V}_{x}$’s decision is:
@@
For the value $\mathfrak{a}$, we have $\mathfrak{a}x^{2}=p_{x}q_{x}{+}p_{x}v_{x}$ and hence $\mathfrak{a}=\frac{1}{y^{2}}(p_{y}q_{y}+p_{y}v_{y})>0$. But then the second derivative $f^{\prime\prime}(x)=2\mathfrak{a}$ of $f$ is positive, which yields that $f$ has a global minimum at some point $x_{0}$ and is strictly decreasing for $x<x_{0}$ and strictly increasing for $x>x_{0}$. As $f(0)$ and $f(1)$ are both negative, we obtain $f(x)<0$ for all $x$ in the interval $[0,1]$. But this contradicts $f(y)\geqslant 0$.
@@
Now, we first provide a finite-memory scheduler $\mathfrak{T}$ for $\mathcal{M}^{\prime}$ that leaves each MEC $\mathcal{E}$ for which $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$ via the state action pair $(s,\alpha)$ with probability $\mathfrak{S}(s_{\mathcal{E}})(\alpha)$. Recall that we assume that each action is enabled in at most one state and that the actions enabled in the state $s_{\mathcal{E}}$ in $\mathit{MEC}(\mathcal{M})$ are precisely the actions that are enabled in some state of $\mathcal{E}$ and that do not belong to $\mathcal{E}$ (Section 2.3)

### v6

@@ 4. Checking the existence of PR causes and the PR conditions
We now turn to algorithms for checking whether a given set $\mathsf{Cause}$ is an SPR or GPR cause for $\mathsf{Eff}$ in $\mathcal{M}$. Since the minimality condition (for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$) of PR causes is verifiable by standard model checking techniques in polynomial time, we concentrate on checking the probability-raising conditions (S) and (G). In the special case where $\mathcal{M}$ is a Markov chain, both conditions (SPR) and (GPR) can be checked in polynomial time by computing the corresponding probabilities. Thus, the interesting case is checking the PR conditions in MDPs. In case of SPR causality this is closely related to the existence of PR causes and decidable in polynomial time (Section 4.1), while checking the GPR condition is more complex and polynomially reducible to (the non-solvability of) a quadratic constraint system (Section 4.2).
We start with the preliminary consideration that for both conditions (S) and (G), it suffices to consider a class of worst-case schedulers, which are minimizing the probability to reach an effect state from every cause state. For this we transform the MDP in question according to the cause candidate in question.
@@
As an example for the model transformation consider the abstract MDP $\mathcal{M}$ from Figure 8 for the singleton set $C=\{c\}$. The transformed MDP $\mathcal{M}_{[c]}$ is seen in Figure 8, where a fresh state $\mathsf{noeff}$ is added. Furthermore, all outgoing transitions from $c$ are deleted a replaced by a fresh action $\gamma$ with exactly two transitions corresponding to $P_{\mathcal{M}_{[c]}}(c,\gamma,\mathsf{eff})=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$ and $P_{\mathcal{M}_{[c]}}(c,\gamma,\mathsf{noeff})=1{-}\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$.
@@
Using $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff})=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\psi_{c}\wedge\Diamond\mathsf{Eff})\ +\ \mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}((\neg\psi_{c})\wedge\Diamond\mathsf{Eff}),$ we can equivalently convert condition (3) for $c$ and $\mathfrak{S}$ to
@@
Suppose $\mathsf{Cause}$ is a GPR cause for $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$. Then, $\mathsf{Cause}$ is a GPR cause for $\mathsf{Eff}$ in $\mathcal{M}$.
@@
From the assumption that $\mathsf{Cause}$ is a GPR cause for $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$, we can conclude that the GPR condition (GPR) holds for all schedulers $\mathfrak{S}$ that satisfy
@@
Pick an MD-scheduler $\mathfrak{T}$ that minimizes the probability to reach $\mathsf{Eff}$ from every state. In particular, $w_{c}=w_{\pi}^{\mathfrak{T}}\leqslant w_{\pi}^{\mathfrak{S}}$ for every state $c\in\mathsf{Cause}$ and every path $\pi\in\Pi_{c}$ (recall that $w_{c}=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$). Moreover, the scheduler $\mathfrak{S}$ can be transformed into a scheduler $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0,\min}$ that is “equivalent” to $\mathfrak{S}$ with respect to the global probability-raising condition. More concretely, let $\mathfrak{S}_{\mathfrak{T}}$ denote the scheduler that behaves as $\mathfrak{S}$ as long as $\mathfrak{S}$ has not yet visited a state in $\mathsf{Cause}$ and behaves as $\mathfrak{T}$ as soon as a state in $\mathsf{Cause}$ has been reached. Thus, $p_{\pi}^{\mathfrak{S}}=p_{\pi}^{\mathfrak{S}_{\mathfrak{T}}}$ and $\mathit{res}(\mathfrak{S}_{\mathfrak{T}},\pi)=\mathfrak{T}$ for each $\pi\in\Pi_{c}$. This yields that the probability to reach $c\in\mathsf{Cause}$ from $\mathsf{init}$ is the same under $\mathfrak{S}$ and $\mathfrak{S}_{\mathfrak{T}}$, i.e., $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond c)=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\Diamond c)$. Therefore $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Cause})=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\Diamond\mathsf{Cause})$. The latter implies that $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0}$, and hence $\mathfrak{S}_{\mathfrak{T}}\in\Sigma_{>0,\min}$. Moreover, $\mathfrak{S}$ and $\mathfrak{S}_{\mathfrak{T}}$ reach $\mathsf{Eff}$ without visiting $\mathsf{Cause}$ with the same probability, i.e., $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,\mathsf{Eff})=\mathrm{Pr}^{\mathfrak{S}_{\mathfrak{T}}}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,\mathsf{Eff})$. But this yields: if (5) holds for $\mathfrak{S}_{\mathfrak{T}}$ then (5) holds for $\mathfrak{S}$. As (5) holds for $\mathfrak{S}_{\mathfrak{T}}$ by assumption, this completes the proof.
@@ 4.1. Checking the strict probability-raising condition and the existence of causes
First, we show the soundness of Algorithm 4.1. By the virtue of Lemma 10 stating the soundness of the transformation $\mathcal{M}$ to $\mathcal{M}_{[c]}$ it suffices to show that Algorithm 4.1 returns the correct answers “yes” or “no” when the task is to check whether the singleton $\mathsf{Cause}=\{c\}$ is an SPR cause in $\mathcal{N}=\mathcal{M}_{[c]}$. Recall the notation $q_{s}=\mathrm{Pr}^{\max}_{\mathcal{M}_{[c]},s}(\Diamond\mathsf{Eff})$. We abbreviate $q=q_{\mathsf{init}}$. Note that $(\neg\mathsf{Cause})\,\mathrm{U}\,c$ is equivalent to $\Diamond c$ as $c\in\mathsf{Cause}$.
@@
    of Algorithm 4.1 (i.e., if $q<w_{c}$), the answer “yes” is sound as $\mathrm{Pr}^{\max}_{\mathcal{N}}(\Diamond\mathsf{Eff})=q<w_{c}$.
@@
    Pick an MD-scheduler $\mathfrak{S}$ of $\mathcal{M}^{\max}_{[c]}$ such that $c$ is reachable from $\mathsf{init}$ via $\mathfrak{S}$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N},s}(\Diamond\mathsf{Eff})=q_{s}$ for all states. Thus, (SPR) does not hold for $c$ and scheduler $\mathfrak{S}$.
@@
    We have $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond c)=0$ for all schedulers $\mathfrak{S}$ for $\mathcal{N}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})=q=w_{c}$. But then $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond c)>0$ implies $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\Diamond\mathsf{Eff})<w_{c}$ as required in (SPR).
@@
    there is a state $c_{0}\in S\setminus\mathsf{Eff}$ such that the singleton $\{c_{0}\}$ is an SPR cause (and therefore a GPR cause) for $\mathsf{Eff}$ in $\mathcal{M}$.
@@ 4.2. Checking the global probability-raising condition
The terminal states $\mathsf{eff}_{\mathsf{unc}}$, $\mathsf{eff}_{\mathsf{cov}}$, $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{noeff}_{\mathsf{tn}}$ are pairwise distinct. $\mathcal{M}$ can have further terminal states representing true negatives. However, these can be identified with $\mathsf{noeff}_{\mathsf{tn}}$.
@@
We justify the assumptions as we can transform $\mathcal{M}$ into a new MDP of the same asymptotic size satisfying the above assumptions. Thanks to Lemma 10, we may suppose that $\mathcal{M}=\mathcal{M}_{[\mathsf{Cause}]}$ without changing the satisfaction of (G). Thus, from cause states $c\in\mathsf{Cause}$ there are only two outgoing transitions, either to a terminal effect state $\mathsf{eff}$ with probability $\mathrm{Pr}^{\min}_{c}(\lozenge\mathsf{Eff})$ or to a terminal non-effect state $\mathsf{noeff}$ with the remaining probability (see Notation 4). We then may rename the effect state $\mathsf{eff}$ and the non-effect state $\mathsf{noeff}$ reachable from $\mathsf{Cause}$ into $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$, respectively. Furthermore, we collapse all other effect states into a single state $\mathsf{eff}_{\mathsf{unc}}$ and all true negative states into a single state $\mathsf{noeff}_{\mathsf{tn}}$. Similarly, by renaming and possibly duplicating terminal states we also suppose that $\mathsf{noeff}_{\mathsf{fp}}$ has no other incoming transitions than the $\gamma$-transitions from the states in $\mathsf{Cause}$. This ensures (A1) and (A2). For (A3) consider the set $T$ of terminal states in the MDP obtained so far. We remove all non-trivial end components by switching to the MEC-quotient [dA97], i.e., we collapse all states that belong to the same MEC $\mathcal{E}$ into a single state $s_{\mathcal{E}}$ representing the MEC while ignoring the actions inside $\mathcal{E}$. Additionally, we add a fresh deterministic $\tau$-transition from the states $s_{\mathcal{E}}$ representing MECs to $\mathsf{noeff}_{\mathsf{tn}}$ (i.e., $P(s_{\mathcal{E}},\tau,\mathsf{noeff}_{\mathsf{tn}})=1$). The $\tau$-transitions from states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ can then be used to mimic cases where the scheduler of the original MDP enters the end component $\mathcal{E}$ and stays there forever.
@@
We demonstrate these transformations on the abstract MDP $\mathcal{M}$ from Figure 12, where the dotted circles correspond to sets of states in the MDP. The MDP already satisfies $\mathcal{M}=\mathcal{M}_{[c]}$. We rename $\mathsf{eff}$ reachable from $\mathsf{Cause}$ to $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}$ to $\mathsf{noeff}_{\mathsf{fp}}$. Effect states not reachable from $\mathsf{Cause}$ collapse to $\mathsf{eff}_{\mathsf{unc}}$. There are no terminal non-effect states not reachable from $c$, which would collapse to $\mathsf{noeff}_{\mathsf{tn}}$. The MEC quotient collapses MECs to states $s_{\mathcal{E}_{i}}$ only keeping outgoing transitions. There is a fresh action $\tau$ in states $s_{\mathcal{E}_{i}}$ to $\mathsf{noeff}_{\mathsf{tn}}$. Thus, we get $\mathcal{N}$ from Figure 12.
@@
Figure 11. MDP $\mathcal{M}$ not satisfying
@@
The violation of (GPR-1) in Lemma 21 and the condition $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ can be reformulated in terms of the frequency-variables as follows where $x_{\mathsf{Cause}}$ is an abbreviation for $\sum_{c\in\mathsf{Cause}}x_{c}$:
@@
We can now prove our first main Theorem 17 of this subsection, stating that deciding the GPR condition can be done in $\mathrm{coNP}$.
@@
There are solutions to (S1)-(S3) and (S5) that violate (S4). The space of feasible points for conditions (S1)-(S3) and (S5) is connected. The right hand side of (S4)
@@
where the variables $y$ are as the original variables $x$ after the variables in $X^{\prime}$ have been removed.
Now, we apply the implicit function theorem: Observe that
@@
The model transformation we used for assumptions (A1)-(A3), however, does affect the memory requirements of a violating scheduler. We may further restrict the MR-schedulers necessary to witness non-causality under assumptions (A1)-(A3). For the following lemma, recall that $\tau$ is the action of the MEC quotient used for the extra transition from states representing MECs to a new trap state (see also assumption (A3)).
@@
To explain how to derive $\mathfrak{T}_{i}$ from $\mathfrak{T}_{i-1}$, let $i\in\{1,\ldots,k\}$, $\mathfrak{V}=\mathfrak{T}_{i-1}$, $u=u_{i}$ and $y=1{-}\mathfrak{V}(u)(\tau)$. By definition we have, $0<y<1$ (as $u\in U$ and by definition of $U$) and $y=\sum_{\alpha\in\mathit{Act}(u)\setminus\{\tau\}}\mathfrak{V}(u)(\alpha)$. For $x\in[0,1]$, let $\mathfrak{V}_{x}$ denote the MR-scheduler that agrees with $\mathfrak{V}$ for all states but $u$, for which $\mathfrak{V}_{x}$’s decision is:
@@
For the value $\mathfrak{a}$, we have $\mathfrak{a}x^{2}=p_{x}q_{x}{+}p_{x}v_{x}$ and hence $\mathfrak{a}=\frac{1}{y^{2}}(p_{y}q_{y}+p_{y}v_{y})>0$. But then the second derivative $f^{\prime\prime}(x)=2\mathfrak{a}$ of $f$ is positive, which yields that $f$ has a global minimum at some point $x_{0}$ and is strictly decreasing for $x<x_{0}$ and strictly increasing for $x>x_{0}$. Since $f(0)$ and $f(1)$ are both negative, we obtain $f(x)<0$ for all $x$ in the interval $[0,1]$. But this contradicts $f(y)\geqslant 0$.
@@
Now, we first provide a finite-memory scheduler $\mathfrak{T}$ for $\mathcal{M}^{\prime}$ that leaves each MEC $\mathcal{E}$ for which $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$ via the state action pair $(s,\alpha)$ with probability $\mathfrak{S}(s_{\mathcal{E}})(\alpha)$. Recall that we assume that each action is enabled in at most one state and that the actions enabled in the state $s_{\mathcal{E}}$ in $\mathit{MEC}(\mathcal{M})$ are precisely the actions that are enabled in some state of $\mathcal{E}$ and that do not belong to $\mathcal{E}$ (Section 2.3).